OpenStudy (loser66):

Suppose that 2000 points are selected randomly and independently from the unit square {(x,y)| 0<=x<1, 0<=y<1} Let W equal number of points that fall into A ={(x,y)|x^2+y^2<1} How W is distributed? Please, help

OpenStudy (loser66):

 to me, n =2000, but how to define p?

OpenStudy (loser66):

Bernoulli distribute need n, p to define W(n, p)

OpenStudy (loser66):

The answer from the book said p =pi/4, but I don't see the logic on it.

OpenStudy (loser66):

 this is the whole thing we have, and A occupied > 1/2 of the square whose sides are 1. How P(A) =pi/4?

OpenStudy (loser66):

@kirbykirby

OpenStudy (kirbykirby):

If A is the region for your event and R is the 2-D region, then \[P(A \in R)=\frac{area ~of ~A}{area~of~R}=\frac{(\pi (1)^2)/4}{1\times 1}\] since the unit square area is 1, and the area of the sector of the circle is 1/4 of the area of the full circle

OpenStudy (loser66):

Oh, I didn't study this formula yet :)

OpenStudy (loser66):

Thanks for the help. May be I will get this formula today.:)

OpenStudy (kirbykirby):

ya it works when you have a uniform distribution over some 2-D region